OpenStudy (anonymous):

given the slope of a line is 7 and it goes through the point (-3, -10) what is the y-intercept???

OpenStudy (anonymous):

well you can use the point slope formula y-y1=m(x-x1) y-(-10)=7(x-(-3)) y+10=7(x+3) y+10=7x+21 -10 -10 y=7x+11 and the y-intercept is the last number in the equation so in this case it would be 11